# VerificAgent: Integrating Expert Knowledge and Fact-Checked Memory for Robust Domain-Specific Task Planning

### v1

@@ Abstract
Continual memory augmentation allows computer-use agents (CUAs) to learn from past interactions and refine their task-solving strategies over time. However, unchecked memory accumulation can introduce spurious or hallucinated “learnings” that degrade agent performance—particularly in domain-specific workflows such as productivity software. We present a novel framework, VerificAgent, that effectively manages memory for CUAs through (1) an expert-curated seed of domain knowledge, (2) iterative, trajectory-based memory refinement during training, and (3) a post-hoc fact-checking pass by human experts to sanitize accumulated memory before deployment. On OSWorld productivity tasks, VerificAgent achieves a 111.1% relative improvement in success rate over baseline CUA without any additional fine-tuning.
footnotetext: † Equal contribution.
@@ 1 Introduction
Autonomous computer-using agents (CUAs) have the potential to transform how users interact with software by delegating routine digital tasks to AI. Recent advances in large language models (LLMs) with multimodal capabilities have enabled agents that can interpret and manipulate graphical user interfaces (GUIs) much like a human operator. For example, OpenAI’s Operator system is trained to read screen content and perform clicks, typing, and other GUI actions without relying on application-specific APIs [1]. This generality allows a single model to work across web and desktop applications. However, while such agents demonstrate impressive versatility, they still struggle with complex, domain-specific tasks. In particular, office productivity applications (e.g., Microsoft Word and PowerPoint) pose a significant challenge for current CUAs, which have been primarily trained on web browsing and simple tool use rather than the intricacies of Office software.
@@
Benchmark evaluations underscore this gap. OSWorld [2] provides a realistic operating system environment with tasks spanning real-world apps, and reports that state-of-the-art agents solve only about 12% of tasks overall, far below human performance ($\approx 72$%). Even with subsequent improvements—OpenAI’s CUA model reached 38% success on OSWorld’s full task suite—performance on Office-specific tasks remains around 25% success in the best models [3, 1]. In practice, current agents often fail at fine-grained Office operations like precise text formatting or slide manipulation, which require remembering complex instructions and adapting to detailed interface layouts. This motivates the need for methods that can boost the task efficiency of CUAs in such complex domains.
@@
Recent work has shown that augmenting LLM agents with persistent memory can substantially improve multi‐step task adaptation and planning [4, 5]. However, these methods do not guard against the gradual accumulation of spurious or hallucinated entries, which can mislead downstream decision‐making [6].
@@
From a theoretical standpoint, verification is often easier than generation—checking a candidate solution requires fewer resources than constructing one—and this distinction carries over to LLMs: empirical studies find that models verify or critique outputs more reliably than they generate them from scratch [7, 8]. Practical frameworks like LLM-Cite demonstrate that fact verification can be performed cheaply and efficiently by generating citation URLs and then validating claims against retrieved documents, reducing verification cost by over an order of magnitude compared to full generative pipelines [9]. Similarly, tools such as SymGen provide direct, in‐context citations that speed human validation of model outputs by presenting source evidence alongside responses [10]. Moreover, in domains where no structured knowledge base exists, a lightweight human review of distilled memory entries can still be carried out with minimal effort, ensuring reliability without major overhead.
Despite these advances in rapid verification, memory‐based agents have largely overlooked the need to sanitize their learned memory stores. No prior system integrates a post-hoc fact-checking pass to prune or correct hallucinated memory entries before inference in a production environment, motivating our practical framework, VerificAgent, which combines expert‐seeded priors, iterative memory learning, and a targeted fact-checking stage to ensure robust, production-ready performance.
@@
  • Scalable Continual Learning Framework. We propose a CUA training pipeline that combines a small, expert-curated domain knowledge base with iterative, trajectory-driven memory updates—allowing the agent to refine its planning strategy over successive trials.
  • Practical Memory Fact-Checking. We introduce a post-hoc verification step in which human experts review and correct the accumulated memory, removing hallucinated or incorrect entries to produce a robust, production-ready knowledge store.
  • Precision Office Grader Enhancements. We improve the OSWorld Office task grader by tightening comparison logic and addressing structural edge cases in PPTX files, yielding more accurate, function-level assessments of agent success.
@@ 2.1 Computer-Using Agents and Benchmarks
OpenAI introduced the Computer-Using Agent (CUA) paradigm with Operator, achieving a 38% success rate on the OSWorld benchmark [1]. OSWorld provides 369 real-world tasks—spanning web browsing, file management, and desktop application use—and highlights a large gap between human ($\approx 72$%) and agent performance (12 %) [2]. OfficeBench zeroes in on multi-step Office productivity workflows, reporting that even GPT-4-based agents plateau at around 47% success [3]. Additional execution benchmarks include DiGIRL for in-the-wild device control [11] and AndroidWorld for mobile GUI agents [12], both underscoring challenges in generalization and error recovery.
@@ 2.2 GUI Grounding and Office Task Automation
Vision-based grounding and screen parsing enable platform-agnostic GUI interaction. OmniParser converts screenshots into structured GUI elements to enhance LLM grounding, boosting performance on ScreenSpot and WebShop benchmarks [13]. UGround trains a universal visual-grounding model on a 10 M-element GUI dataset, supporting pixel-level operations across desktop, web, and mobile without relying on accessibility APIs [14]. WinClick advances Windows-specific grounding by combining screenshot-based detection with LLM alignment, achieving superior results on the WinSpot benchmark [15]. Unlike these perception-focused systems, our work targets the decision-making layer by integrating initial domain knowledge injection, trajectory-based memory, and post-hoc fact-checking to drive robust planning for Office-specific tasks.
@@ 2.3 Compositional and Modular Agent Architectures
To tackle the complexity of long-horizon GUI tasks, recent works propose hierarchical and modular designs. Agent S introduces an experience-augmented hierarchical planner and an Agent–Computer Interface (ACI) for domain knowledge acquisition and dynamic interface handling, setting new SOTA on OSWorld and WindowsAgentArena [16]. Agent S2 extends this with a Compositional Generalist–Specialist framework—using a Mixture-of-Grounding mechanism for precise GUI localization and Proactive Hierarchical Planning—outperforming monolithic agents across OSWorld, WindowsAgentArena, and AndroidWorld benchmarks [17]. UI-TARS pioneers a native end-to-end GUI agent that learns from iterative reflective traces and vision grounding, achieving 24.6% success on OSWorld (50-step) and 46.6% on AndroidWorld using only screenshots as input [18].
@@ 2.4.1 Planning
Interleaving reasoning with action has become a cornerstone of LLM‐driven agents: ReAct combines chain‐of‐thought traces with environment actions to reduce hallucinations in single episodes [19], and Reflexion adds a self‐reflection loop that stores textual “reflections” as episodic memory for future planning [20]. Retrieval‐Augmented Generation (RAG) further boosts plan reliability by injecting external context—surveys of RAG for meeting‐scale LLMs [21] and contextual memory mechanisms like RAP [22] demonstrate substantial gains. Tool‐use frameworks such as Toolformer [23], HuggingGPT [24], and Gorilla [25] enable agents to invoke APIs and external tools within their plans, while chain‐based systems like Auto‐GPT orchestrate LLM‐generated plans and tool calls toward complex goals, often augmented by external memory stores for long‐horizon coordination [26]. In web navigation, RADA integrates search‐based retrieval into multi‐step planning [27], and the AutoGuide series applies state‐aware guideline selection with retrieval‐augmented prompts to curb hallucinations in web‐driven interactions [28, 29].
While prior work demonstrates the power of planning, it rarely examines how the planner’s inputs—especially varying levels of domain knowledge—affect plan quality and success. In this paper, we perform a deeper analysis of planning under three knowledge regimes (expert seed, raw memory, and fact‐checked memory), quantifying how each influences the accuracy, efficiency, and robustness of generated GUI action sequences.
@@ 2.4.2 Memory
Recent LLM agents have demonstrated the value of explicit memory modules for capturing and reusing experience over multiple trials. For instance, CLIN maintains a dynamic cache of causal abstractions to enable rapid adaptation without weight updates [4], LongMem pairs a frozen backbone LLM with a residual side‐network for long‐term context retrieval [5], and TiM employs iterative recall and post‐thinking stages to evolve session‐level “thoughts” [6]. A comprehensive survey further categorizes these and other memory designs—covering object, episodic, and procedural patterns—across diverse applications [30]. Retrieval‐augmented methods (e.g. RAG) inject external knowledge to reduce hallucinations in generated outputs [21], but they do not address errors that accumulate within the memory store itself.
To our knowledge, no prior system incorporates a dedicated fact‐checking stage to prune or correct spurious memory entries before deployment. We present the first practical memory sanitization framework—combining human‐in‐the‐loop validation with lightweight automated checks—to eliminate hallucinated learnings and make dynamic memory safe and reliable for production‐ready GUI agents.
3 Method
@@ 3.1 Problem Formulation
[…]
@@
    The agent is given the natural‐language task description $\mathcal{T}$ and an initial screenshot $S_{0}$ capturing the current slide state.
@@
    After executing the final action $a_{k}$, the agent downloads the resulting PowerPoint deck $D_{k}$.
@@
$g\bigl{(}T,D_{k}\bigr{)}=\begin{cases}1,&\text{if }D_{k}\text{satisfies}\mathcal{T},\\ 0,&\text{otherwise.}\end{cases}$
@@
(a) Learning workflow At each iteration, the Task Generator dispenses a new task $\mathcal{T}$ with an initial screenshot $\mathcal{S}_{0}$. Expert‐curated domain knowledge $\mathcal{K}$, the current memory $\mathcal{M}$, and the task ($\mathcal{T},\mathcal{S}_{0}$) seed the Plan Generator, which emits a step‐by‐step plan $\mathcal{P}$. The Computer‐Use Agent executes $\mathcal{P}$ in the developer environment, producing a saved trajectory $\mathcal{\tau}$. A grader analyzes the outcome and outputs a score $g$. The Analyzer ingests $\bigl{(}\mathcal{\tau},\,\mathcal{P},g\bigr{)}$ and updates $\mathcal{M}$, closing the loop and informing the next planning cycle.
(b) Inference workflow After training, the learned memory $\mathcal{M}$ is manually fact-checked by human experts to remove any erroneous entries, yielding a frozen memory $\mathcal{M}^{*}$. During inference, the Planner takes the new task–screenshot pair $(\mathcal{T},\mathcal{S}_{0})$ and $\mathcal{M}^{*}$ as inputs to generate a step-by-step plan $\mathcal{P}$, which the Computer-Use Agent then executes in the production environment. No further memory updates occur during inference.
@@
3.2 Overall System Architecture
Our framework (Figure 1) operates in two stages: a learning phase, during which memory is iteratively built and refined, and an inference phase, in which the expert–fact-checked memory is frozen and used—without further updates—to execute tasks on the test set.

### v2

# VerificAgent: Domain-Specific Memory Verification for Scalable Oversight of Aligned Computer-Use Agents

@@ Abstract
Continual memory augmentation lets computer-using agents (CUAs) learn from prior interactions, but unvetted memories can encode domain-inappropriate or unsafe heuristics–spurious “rules” that drift from user intent and safety constraints. We introduce VerificAgent, a scalable oversight framework that treats persistent memory as an explicit alignment surface. VerificAgent combines (1) an expert-curated seed of domain knowledge, (2) iterative, trajectory-based memory growth during training, and (3) a post-hoc human fact-checking pass to sanitize accumulated memories before deployment. Evaluated on OSWorld productivity tasks and additional adversarial stress tests, VerificAgent improves task reliability, reduces hallucination-induced failures, and preserves interpretable, auditable guidance–without additional model fine-tuning. By letting humans correct high-impact errors once, the verified memory acts as a frozen safety contract that future agent actions must satisfy. Our results suggest that domain-scoped, human-verified memory offers a scalable oversight mechanism for CUAs, complementing broader alignment strategies by limiting silent policy drift and anchoring agent behavior to the norms and safety constraints of the target domain.
@@ 1 Introduction
Autonomous computer-using agents (CUAs) have the potential to transform how users interact with software by delegating routine digital tasks to AI. Recent advances in large language models (LLMs) with multimodal capabilities have enabled agents that can interpret and manipulate graphical user interfaces (GUIs) much like a human operator. For example, OpenAI’s Operator system is trained to read screen content and perform clicks, typing, and other GUI actions without relying on application-specific APIs openai_operator2025 . This generality allows a single model to work across web and desktop applications. However, while such agents demonstrate impressive versatility, they still struggle with complex, domain-specific tasks. In particular, office productivity applications (e.g., Microsoft Word and PowerPoint) pose a significant challenge for current CUAs, which have been primarily trained on web browsing and simple tool use rather than the intricacies of Office software.
@@
Benchmark evaluations underscore this gap. OSWorld Xie2024OSWorld provides a realistic operating system environment with tasks spanning real-world apps, and reports that state-of-the-art agents solve only about 12% of tasks overall, far below human performance (72$\approx$ 72≈ 72%). Even with subsequent improvements—OpenAI’s CUA model reached 38% success on OSWorld’s full task suite—performance on Office-specific tasks remains around 25% success in the best models Wang2024OfficeBench ; openai_operator2025 . In practice, current agents often fail at fine-grained Office operations like precise text formatting or slide manipulation, which require remembering complex instructions and adapting to detailed interface layouts. This motivates the need for methods that can boost the task efficiency of CUAs in such complex domains.
@@
Recent work has shown that augmenting LLM agents with persistent memory can substantially improve multi-step task adaptation and planning Majumder2023CLIN ; Wang2023LongMem . However, these methods do not guard against the gradual accumulation of spurious or hallucinated entries, which can mislead downstream decision-making Liu2023TiM . Unvetted memories can encode latent policies that violate user intent or safety constraints, so verification becomes a form of scalable oversight.
@@
From a theoretical standpoint, verification is often easier than generation—checking a candidate solution requires fewer resources than constructing one—and this distinction carries over to LLMs: empirical studies find that models verify or critique outputs more reliably than they generate them from scratch Stechly2024SelfVerification ; Liang2023BetterReasoners . Practical frameworks like LLM-Cite demonstrate that fact verification can be performed cheaply and efficiently by generating citation URLs and then validating claims against retrieved documents, reducing verification cost by over an order of magnitude compared to full generative pipelines Qiu2024LLMCite . Similarly, tools such as SymGen provide direct, in-context citations that speed human validation of model outputs by presenting source evidence alongside responses MitCSAIl2024Verify . Moreover, in domains where no structured knowledge base exists, a lightweight human review of distilled memory entries can still be carried out with minimal effort, ensuring reliability without major overhead.
Despite these advances in rapid verification, memory-based agents have largely overlooked the need to sanitize their learned memory stores. No prior system integrates a post-hoc fact-checking pass to prune or correct hallucinated memory entries before inference in a production environment. From an alignment perspective, memory-augmented agents face a distinct risk: they may internalize hallucinated ‘policies’ that deviate from user values and safety constraints. Our central question is how to ensure that an agent’s persistent memory remains accurate and value-consistent as capabilities scale. We treat memory verification as a form of scalable oversight: humans correct high-impact errors once, then freeze these corrections to guide future behaviors.
In this work, we cast persistent agent memory as both a capability enabler and a latent alignment threat: without safeguards, agents may internalize spurious rules that induce unsafe or incorrect behaviors. VerificAgent addresses this by combining expert-seeded priors, iterative memory growth, and a human-in-the-loop sanitization step, yielding a scalable oversight pipeline that aligns the agent’s learned heuristics with domain-specific safety constraints.
@@
  1. 1.
    VerificAgent: scalable oversight that boosts capability and safety. We introduce a two-stage “generate-then-freeze” pipeline that turns persistent memory from an alignment liability into an asset—nearly doubling task success on the OSWorld Office benchmark while cutting harmful actions, without any extra model training.
  2. 2.
    Alignment-centric evaluation suite for GUI agents. VerificAgent is integrated into two public agents—OpenAI computer-use-preview and ByteDance UI-TARS-1.5-7B—and assessed on both the OSWorld Office benchmark and a new Red-Team Memory Injection testbed, providing the community with a reproducible yardstick for oversight methods.
  3. 3.
    Generalizable mitigation blueprint for the “unfiltered-memory” alignment gap. Our analysis formalizes a systemic risk shared by memory-augmented agents and distills a domain-agnostic checklist—multi-layered memory-update, AI assisted oversight, and domain expert approval—to guide safe deployment in other high-stakes domains.
@@ 2.1 Computer-Using Agents and Benchmarks
OpenAI introduced the CUA (CUA) paradigm with Operator, achieving a 38% success rate on the OSWorld benchmark openai_operator2025 . OSWorld provides 369 real-world tasks—spanning web browsing, file management, and desktop application use—and highlights a large gap between human (72$\approx$ 72≈ 72%) and agent performance (12 %) Xie2024OSWorld . OfficeBench zeroes in on multi-step Office productivity workflows, reporting that even GPT-4-based agents plateau at around 47% success Wang2024OfficeBench . Additional execution benchmarks include DiGIRL for in-the-wild device control Bai2024DiGIRL and AndroidWorld for mobile GUI agents Rawles2024AndroidWorld , both underscoring challenges in generalization and error recovery.
@@ 2.2 GUI Grounding and Office Task Automation
Vision-based grounding and screen parsing enable platform-agnostic GUI interaction. OmniParser converts screenshots into structured GUI elements to enhance LLM grounding, boosting performance on ScreenSpot and WebShop benchmarks Lu2024OmniParser . UGround trains a universal visual-grounding model on a 10 M-element GUI dataset, supporting pixel-level operations across desktop, web, and mobile without relying on accessibility APIs Gou2024UGround . WinClick advances Windows-specific grounding by combining screenshot-based detection with LLM alignment, achieving superior results on the WinSpot benchmark Hui2025WinClick . Unlike these perception-focused systems, our work targets the decision-making layer by integrating initial domain knowledge injection, trajectory-based memory, and post-hoc fact-checking to drive robust planning for Office-specific tasks.
@@ 2.3 Compositional and Modular Agent Architectures
To tackle the complexity of long-horizon GUI tasks, recent works propose hierarchical and modular designs. Agent S introduces an experience-augmented hierarchical planner and an Agent–Computer Interface (ACI) for domain knowledge acquisition and dynamic interface handling, setting new SOTA on OSWorld and WindowsAgentArena Agashe2024AgentS . Agent S2 extends this with a Compositional Generalist–Specialist framework—using a Mixture-of-Grounding mechanism for precise GUI localization and Proactive Hierarchical Planning—outperforming monolithic agents across OSWorld, WindowsAgentArena, and AndroidWorld benchmarks Agashe2025AgentS2 . UI-TARS pioneers a native end-to-end GUI agent that learns from iterative reflective traces and vision grounding, achieving 24.6% success on OSWorld (50-step) and 46.6% on AndroidWorld using only screenshots as input Qin2025UITARS .
@@ 2.4.1 Planning
Interleaving reasoning with action has become a cornerstone of LLM-driven agents: ReAct combines chain-of-thought traces with environment actions to reduce hallucinations in single episodes Yao2022ReAct , and Reflexion adds a self-reflection loop that stores textual “reflections” as episodic memory for future planning Shinn2023Reflexion . Retrieval-Augmented Generation (RAG) further boosts plan reliability by injecting external context—surveys of RAG for meeting-scale LLMs Fan2024RAG and contextual memory mechanisms like RAP Kagaya2024RAP demonstrate substantial gains. Tool-use frameworks such as Toolformer Schick2023Toolformer , HuggingGPT Shen2023HuggingGPT , and Gorilla Patil2023Gorilla enable agents to invoke APIs and external tools within their plans, while chain-based systems like Auto-GPT orchestrate LLM-generated plans and tool calls toward complex goals, often augmented by external memory stores for long-horizon coordination Yang2023AutoGPT . In web navigation, RADA integrates search-based retrieval into multi-step planning Kim2024RADA , and the AutoGuide series applies state-aware guideline selection with retrieval-augmented prompts to curb hallucinations in web-driven interactions Fu2024AutoGuideA ; Fu2024AutoGuideB .
While prior work demonstrates the power of planning, it rarely examines how the planner’s inputs—especially varying levels of domain knowledge—affect plan quality and success. In this paper, we perform a deeper analysis of planning under three knowledge regimes (expert seed, raw memory, and fact-checked memory), quantifying how each influences the accuracy, efficiency, and robustness of generated GUI action sequences.
@@ 2.4.2 Memory
Recent LLM agents have demonstrated the value of explicit memory modules for capturing and reusing experience over multiple trials. For instance, CLIN maintains a dynamic cache of causal abstractions to enable rapid adaptation without weight updates Majumder2023CLIN , LongMem pairs a frozen backbone LLM with a residual side-network for long-term context retrieval Wang2023LongMem , and TiM employs iterative recall and post-thinking stages to evolve session-level “thoughts” Liu2023TiM . A comprehensive survey further categorizes these and other memory designs—covering object, episodic, and procedural patterns—across diverse applications Survey2024LLMAgentMemory . Retrieval-augmented methods (e.g. RAG) inject external knowledge to reduce hallucinations in generated outputs Fan2024RAG , but they do not address errors that accumulate within the memory store itself.
2.5 Memory as an Alignment Surface
The dominant line of alignment research frames the problem as learning a reward that captures human preferences, beginning with deep reinforcement learning from human preferences (RLHF) christiano2023deepreinforcementlearninghuman . Scaling this paradigm to large language models showed that supervised + RLHF fine-tuning can elicit instruction-following behavior in systems such as InstructGPT ouyang2022traininglanguagemodelsfollow . To lower the cost of human supervision, recent work replaces or supplements human labels with model-generated feedback—e.g., RL from AI Feedback (RLAIF) and Constitutional AI—which preserves alignment quality while improving scalability lee2024rlaifvsrlhfscaling ; bai2022constitutionalaiharmlessnessai . A complementary thread on scalable oversight amplifies evaluators through recursive reward modeling leike2018scalableagentalignmentreward , debate between competing agents irving2018aisafetydebate ; browncohen2023scalableaisafetydoublyefficient , and other amplification protocols bridgers2024human . Game-theoretic analyses such as the Off-Switch Game emphasize the importance of mechanisms that let humans intervene after deployment rather than solely at training time hadfieldmenell2017offswitchgame . Yet these approaches predominantly act on the policy or reward function, leaving an agent’s persistent memory—where extracted knowledge and habits accumulate—largely unchecked. Contemporary memory-augmented agents record factual corrections and task heuristics without explicit safety guarantees Agashe2024AgentS .
Anthropic’s recent ARC-Evals auditing agents automate whole-model investigations: an investigator uncovers hidden goals, a red-team agent elicits unsafe completions, and an evaluation agent builds bespoke behavioural evals marks2025auditing . While our pipeline shares their goal of scalable oversight, it targets a different alignment surface: instead of probing latent weights, we freeze a vetted memory bank that the agent must query at inference time. This “verify-then-freeze” paradigm turns memory into a hard security boundary and gives downstream auditors a single, human-readable artefact to certify—something ARC-style auditors deliberately avoid to keep models fully intact.
SAFE (Search-Augmented Factuality Evaluator) decomposes generated text into atomic claims and verifies each via web search, cutting long-form fact-checking cost by >20×>20$\times$> 20 × while matching human raters on 72% of claims wei2024longformfactualitylargelanguage . However, SAFE presumes that the requisite ground-truth knowledge is publicly indexed; our setting is domain-specific—enterprise Office workflows—where much of the authoritative knowledge (e.g., custom macros, proprietary templates, internal policy) is not available on the open Web, so web-search verification is infeasible. VerificAgent therefore inverts SAFE’s workflow: we pre-approve knowledge before it can influence generation, avoiding SAFE’s failure mode in which hallucinated content is surfaced to users first and only flagged afterwards. SAFE’s claim-level granularity also makes it less suited to structural threats such as “plan-injection” attacks that manipulate task graphs rather than individual facts.
Chen et al.’s AgentPoison shows how back-door triggers embedded in a web agent’s long-term memory can hijack future plans with $>$ 80% success at $<$ 0.1% poison rate chen2024agentpoison . Our adversarial evaluation corroborates that threat but further demonstrates that a two-tier filter plus human veto eliminates poisoned entries entirely (0% infiltration in RTMI). VerificAgent treat memory as an attack vector, yet we adopt purge-style sanitisation and freeze the vetted subset as an immutable read-only table.
Complementary studies reinforce the need for such defences. MINJA shows fully interactive memory-injection without direct DB access dong2025practicalmemoryinjectionattack ; “plan-injection” corrupts task graphs in browser agents even with prompt-sandboxing patlan2025contextmanipulationattacks ; MemBench exposes how many agents trade off memory capacity for retrieval noise, complicating naive black-box auditing tan2025membenchcomprehensiveevaluationmemory ; and memory back-door attacks on vision nets highlight cross-modality generality of the threat luzon2024memorybackdoorattacksneural . VerificAgent addresses all three failure channels—accidental hallucination, systematic poisoning, and context-manipulation—by making memory state the primary locus of oversight.
3 VerificAgent Framework
@@ 3.1 Problem Formulation
We address the general problem of executing a high-level, domain-specific task $\mathcal{T}$ within an interactive GUI environment—an abstraction that applies to any software application. We instantiate our framework in the web-based Microsoft Office apps, including Word, Excel, and PowerPoint.
@@
    The agent is given the natural-language task description $\mathcal{T}$ and an initial screenshot $S_{0}$ capturing the current slide state.
@@
    After executing the final action $a_{k}$, the agent downloads the resulting Office document $D_{k}$.
@@
$g\bigl{(}T,D_{k}\bigr{)}=\begin{cases}1,&\text{if }D_{k}~\text{satisfies}~\mathcal{T},\\ 0,&\text{otherwise.}\end{cases}$
@@
(a) Learning workflow At each iteration, the Task Generator dispenses a new task $\mathcal{T}$ with an initial screenshot $\mathcal{S}_{0}$. Expert-curated domain knowledge $\mathcal{K}$, the current memory $\mathcal{M}$, and the task ($\mathcal{T},\mathcal{S}_{0}$) seed the Plan Generator, which emits a step-by-step plan $\mathcal{P}$. The Computer-Use Agent executes $\mathcal{P}$ in the developer environment, producing a saved trajectory $\mathcal{\tau}$. A grader analyzes the outcome and outputs a score $g$. The Analyzer ingests $\bigl{(}\mathcal{\tau},\,\mathcal{P},g\bigr{)}$ and updates $\mathcal{M}$, closing the loop and informing the next planning cycle.
(b) Memory oversight After training, the learned memory $\mathcal{M}$ is manually fact-checked by human experts, aided by AI suggestions, to remove any erroneous entries, yielding a frozen memory $\mathcal{M}^{*}$.
(c) Inference workflow During inference, the Planner takes the new task–screenshot pair $(\mathcal{T},\mathcal{S}_{0})$ and $\mathcal{M}^{*}$ as inputs to generate a step-by-step plan $\mathcal{P}$, which the Computer-Using Agent then executes in the production environment. No further memory updates occur during inference.
@@
3.2 System Architecture
Our framework (Figure 1) operates in three stages: a learning phase, during which memory is iteratively built and refined, a memory oversight phase, where domain experts fact-check the learned memory for alignment, and an inference phase, in which the reviewed memory is frozen and used without further updates to execute tasks on the test set.